# Beyond Perception: Evaluating Abstract Visual Reasoning through Multi-Stage Task

### v1

@@ Abstract
Current Multimodal Large Language Models (MLLMs) excel in general visual reasoning but remain underexplored in Abstract Visual Reasoning (AVR), which demands higher-order reasoning to identify abstract rules beyond simple perception. Existing AVR benchmarks focus on single-step reasoning, emphasizing the end result but neglecting the multi-stage nature of reasoning process. Past studies found MLLMs struggle with these benchmarks, but it doesn’t explain how they fail. To address this gap, we introduce MultiStAR, a Multi-Stage AVR benchmark, based on RAVEN, designed to assess reasoning across varying levels of complexity. Additionally, existing metrics like accuracy only focus on the final outcomes while do not account for the correctness of intermediate steps. Therefore, we propose a novel metric, MSEval, which considers the correctness of intermediate steps in addition to the final outcomes. We conduct comprehensive experiments on MultiStAR using 17 representative close-source and open-source MLLMs. The results reveal that while existing MLLMs perform adequately on basic perception tasks, they continue to face challenges in more complex rule detection stages. The dataset and code are available at https://github.com/YanbeiJiang/MultiStAR
@@
Figure 1: Left Part: RAVEN puzzle. The correct answer is 1. Right Part: Direct Answer subtask, where questions are independent for each configuration; Logical Chain subtask, where information from previous stages is used to assist in answering the current stage, all questions here focus on the concept of Number.
@@ 1 Introduction
Multimodal Large Language Models (MLLMs) demonstrate proficiency in addressing a wide array of visual-text inquiries and show strong multimodal understanding ability in tasks such as visual question answering Goyal et al. (2017); Marino et al. (2019); Ding et al. (2023), image captioning Saito et al. (2023); Vinyals et al. (2015); Jiang et al. (2024a), and visual grounding He et al. (2024); Deng et al. (2021). These tasks focus on evaluating the models’ capability to understand real-world or domain-specific knowledge. However, Abstract Visual Reasoning (AVR) presents a different challenge, focusing on a model’s ability to identify and reason through abstract patterns, relationships, and rules. A well-known example of AVR tasks is RAVEN Raven (2003); Zhang et al. (2019), as shown in the left part of Figure 1. The solver needs to select the correct panel from a answer set to complete a 3x3 problem matrix by deducing the visual rules governing the grid’s arrangement. For instance, by analyzing the colors of each panel, one might observe the color remains consistent across each row. Unlike other multimodal tasks in real-world scenarios, AVR focuses on reasoning about arbitrary visual elements, serving as a robust benchmark for evaluating the zero-shot reasoning capabilities of MLLMs in visual contexts Mańdziuk and Żychowski (2019); Santoro et al. (2018).
@@
Previous works have consistently shown that AVR tasks pose challenges for MLLMs in zero-shot inference settings. Despite recent advancements like Chain-of-Thought prompting Ahrabian et al. (2024); Gendron et al. (2024) and the inclusion of oracle captions Zhang et al. (2024), models continue to perform at near-random levels on these tasks. The AVR datasets used commonly in these evaluations like RAVEN primarily focus on single-step end-to-end reasoning (i.e., giving the models the questions and asking them to derive the final answer), as shown in the left part of Figure 1 Santoro et al. (2018); Nie et al. (2020); Cao et al. (2024). However, this design deviates from the human reasoning process, which often involves sequential steps: starting with single-panel perception, progressing to panel comparisons, and finally deducing the underlying rules before solving the puzzle. Previous datasets often omit these intermediate stages, posing challenges to effectively evaluate their step-by-step reasoning capabilities and identify where models struggle within the reasoning process. This highlights the need for benchmarks that assess intermediate perception and reasoning processes. Additionally, a model that accurately identifies patterns in early steps but fails in the final deduction still demonstrates partial reasoning capability. Rewarding such intermediate success aligns with human evaluation practices. However, existing metrics like accuracy, measure only the performance of the current stage while disregarding the correctness of intermediate steps.
@@
To address the limitation of lacking intermediate process evaluation, we introduce MultiStAR, a Multi-Stage Abstract Visual Reasoning dataset, designed to evaluate MLLMs on the intermediate steps in the reasoning process. As shown in Figure 1, the dataset is divided into two sub-tasks, each focusing on different aspects of reasoning. The first sub-task, referred to as Direct Answer, evaluates model performance at varying levels of complexity to assess perception and reasoning abilities at each individual step. Using template-based methods, we generate questions based on RAVEN, ranging from basic object recognition to advanced comparison, pattern recognition, and rule inference. This approach ensures comprehensive coverage of reasoning patterns. The second sub-task called Logical Chain, emphasizes how models measure and maintain logical correlations across reasoning steps. Using puzzles from the RAVEN as the final question, we decompose the reasoning process into a sequence of subproblems in a bottom-up manner. Each stage in this chain links the current reasoning task to its dependent subproblems, requiring the model to combine current information with outputs from previous stages. To assess the correctness of intermediate steps, we introduce a novel metric, MSEval, which provides a more fine-grained assessment of the model’s reasoning process for the logical chain task. MSEval uses the correct answer probabilities at each stage to compute the joint probability across the reasoning process. This approach considers both the correctness of the current stage and all dependent intermediate steps.
@@
In summary, our contributions are: 1) We introduce the MultiStAR benchmark, designed to evaluate models across different stages of reasoning through two subtasks, allowing for a more granular analysis of their performance throughout the reasoning process. 2) We present a novel metric that incorporates the correctness of the current stage as well as the accuracy of its dependent intermediate steps. 3) We perform extensive experiments on a wide range of state-of-the-art MLLMs, providing insights into their strengths, weaknesses, and future improvement directions on AVR tasks.
Figure 2: Left Part: Direct Answer subtask, showcasing six configurations along with their corresponding examples. Right Part: Logical Chain task, presenting a partial view of the logical chain (See full chain and the chain designing rationale in Appendix A.3). Examples are provided for one specific path in the chain.
@@ 2 Related Work
Visual reasoning benchmarks have evolved to assess the capacity of AI models on various tasks including compositional Johnson et al. (2017), commonsense Gao et al. (2022); Li et al. (2024), scientific Hiippala et al. (2021); Saikh et al. (2022); Yue et al. (2024), and abstract visual reasoning. Both commonsense and scientific reasoning tasks require real-world knowledge and a prior understanding of specific domains. Abstract Visual Reasoning (AVR) benchmarks the main focus of this work, primarily involving classification tasks, where models select an answer from a fixed set of choices based on abstract patterns and rules Mańdziuk and Żychowski (2019); Zhang et al. (2019); Santoro et al. (2018); Nie et al. (2020). A few other AVR benchmarks address generative tasks, where models are tasked with recreating elements that fit within a given visual sequence, introducing additional complexity by evaluating a model’s creative reasoning capabilities Chollet (2019); Moskvichev et al. (2023). The most similar benchmark to ours is MARVEL Jiang et al. (2024b), which targets AVR tasks and extends reasoning diversity with six core patterns across geometric and abstract shapes. It also includes basic perception questions to assess visual comprehension. However, MARVEL is still limited in its capacity to analyze intermediate reasoning steps. Table 1 shows the key statistics and features comparison of major multimodal reasoning datasets alongside our proposed MultiStAR benchmark.
@@ 4 Results
In the Direct Answer subtask, we evaluate 17 representative MLLMs in zero-shot setting, including both open-source and close-source models. For open-source models, we have pre-trained (without instruction-tuned) and instruction-tuned models. Additionally, we consider a range of model sizes to ensure a comprehensive evaluation. The model settings and the input prompt details are shown in Appendix A.4.1.
@@ 4.1 Result Analysis
Table 2 compares the performance of various MLLMs on the Direct Answer task. The two close-source models, GPT-4o and Gemini-1.5-pro, outperform others in 1P-B, 1R and 2R stages. GPT-4o achieves an impressive 88.07% accuracy for basic object-oriented questions within a single panel, highlighting its strong capability to recognize simple visual patterns. Gemini achieves the best results in rule deduction tasks for both one-row and two-row configurations, indicating its superior ability to process more complex visual inputs and perform logical reasoning effectively. Among open-source models, pre-trained models generally perform worse than instruction-tuned ones, emphasizing that instruction tuning is an effective approach for addressing question-answering and reasoning tasks. Qwen2-VL-Instruct-72B achieves the best performance on cross-panel comparison tasks, showcasing its strength in identifying relationships between objects across panels. Regarding the limitations of the models, we performed an error analysis (refer to Appendix 11) and identified that the models mainly face challenges with deep reasoning errors. Additionally, the decode-only models, such as Qwen2-VL, InternVL2, NVLM-D, tend to perform much better than encoder-decoder models like InstructBLIP or LLaVA-v1.5. This suggests that decoder-only architectures may be more effective for step-by-step reasoning to some extent. However, the observed performance gap may not be only attributed to architectural differences, there are other potential confounds such as improved training data or additional new functionalities.
Interestingly, as the questions become increasingly complex and require deeper reasoning, a noticeable decline in performance is observed across all models, gradually approaching the random baseline. While models demonstrate strong performance on basic perception tasks, they struggle significantly with deeper reasoning challenges. In contrast, human performance remains stable at above 60% with increasing complexity. This highlights the substantial gap between model and human performance, emphasizing the limitations of current MLLMs in understanding and reasoning at a level comparable to humans.
Another finding is that the performance generally improves with larger model sizes (See Appendix A.9.1 for detailed visual analysis.), mainly due to differences in the size of their language encoders. This indicates that a robust language encoder significantly influences overall performance.
@@
Table 2: The answer accuracy of MLLMs for the Direct Answer subtask. The best results are highlighted in bold. The model may have included the RAVEN dataset in training, these results are no longer comparable as baselines. Result reported in Zhang et al. (2019).
@@
Table 3: The Accuracy (Acc) and MSEval scores for the Logical Chain task. *The model may have included the RAVEN dataset in training, these results are no longer comparable as baselines. The highest accuracy are highlighted in bold. The highest MSEval are highlighted in underline. w/o: without prior, w: with prior. See Appendix A.4.2 for results of more models.
@@
Table 3 shows the performance of MLLMs on each stage of the Logical Chain task. Results are reported as accuracy and MSEval scores. Among all models, Gemini-1.5-pro achieves the best performance on the first four stages, demonstrating its superior ability to reason through multi-stage dependencies. Among the open-source models, Qwen2-VL-72B outperforms others in both accuracy and MSEval, suggesting that our MSEval metric generally aligns with accuracy. When comparing results with and without prior, all models show better performance when prior is available in both metrics, highlighting their capacity to benefit from step-by-step reasoning, even when some generated previous answers might be incorrect. For visual representation of the percentage increase, refer to Appendix A.7.
Interestingly, for the final stage involving the RAVEN puzzle, prior information appears to provide limited utility for most models, with accuracy close to random except for those that might touch on RAVEN tasks. This aligns with previous findings that chain-of-thought reasoning models struggle to solve RAVEN puzzles Ahrabian et al. (2024); Gendron et al. (2024). However, MSEval scores tell a different story. Models with prior information, particularly NVLM-D-72B, show significant improvements in MSEval (close to 100%), despite low final stage accuracy. Since MSEval evaluates correctness across intermediate and current stages, it reveals that models, despite failing the final stage, often solve intermediate steps with higher confidence. Another finding is the MSEval score for Qwen2-VL-72B declines when provided with prior information, which is not consistent with accuracy. This indicates its weakness in addressing intermediate stages that were likely not part of its pre-training. Despite errors in earlier stages, the model still performs well on the final question, suggesting it relies on memorizing patterns from the final stage rather than demonstrating a strong understanding of the logical reasoning behind the task. This highlights a critical limitation in current MLLMs, while they may achieve impressive results in isolated cases, their ability to generalize and reason through multi-stage logical dependencies remains inadequate. To further verify the MSEval’s effectiveness, we also conduct qualitative analysis in section 6.
@@ 5 Discussion
From Figure 5, the “number” attribute is the easiest to recognize in higher level configurations (2P, 1R, 2R), while “position” is the most easily identified in low-level, single-panel settings. Some attributes achieve accuracy above 90%, indicating that the models exhibit strong counting and spatial reasoning capabilities. However, they struggle with attributes like “color” and “size”, particularly in high-level configurations, suggesting that the models may not be adequately designed or trained to focus on these aspects.
Figure 5: Breakdown analysis of five attributes for Gemeni-1.5-pro and Qwen2-VL-72B on the Direct Answer task. Refer to the Appendix A.9.2 for more models.
Given the ground truth for intermediate steps, how does it influence the final results?
Table 4 highlights the ground truth priors generally demonstrate a positive impact. For example, the 1R stage benefits significantly from the insights about each panel and intra-panel comparisons. The 2R stage also sees substantial gains, as it mainly relies on double-checking information from the 1R stage without requiring additional changes in most cases. However, the final stage experiences a negative impact despite the inclusion of correct rules. This may be attributed to the complexity of the visual input, which contains numerous objects, making it challenging for the model to effectively apply the given rules. And for the Qwen2-VL-72B model, its tendency to memorize patterns might turn these ground truths into noise.
Prior Info 1P 2P 1R 2R Final
@@
Table 4: The accuracy with incorporating ground truth information at each stage for Logical Chain task. *These results are no longer comparable as baselines.
@@
Injecting prior information into prompts significantly increases their length (see Appendix A.1 for details), making it more challenging for models to focus on critical details. To address this issue, we proposed two methods: (1) adding HTML tags to structure the prompt by separating prior information, background, and questions, enabling the model to clearly distinguish each part, and (2) formatting the prompt as a PDF document with distinct sections and titles. Table 6 demonstrates that HTML tagging provides notable benefits, particularly Qwen2-VL and GPT-4o, while the document-based approach proves less effective, especially for high-level stages. However, for other open-source models, they yields no improvements (see Appendix A.9.3 for further details and examples of the conversion methods).
Prior 1P 2P 1R 2R Final
@@
Table 6: The accuracy of three prompting techniques for prior information for Logical Chain task. Vanilla: Pure Text, Struct.: Structure (HTML), Doc.: Document. *These results are no longer comparable as baselines.

### v2

@@ Abstract
Current Multimodal Large Language Models (MLLMs) excel in general visual reasoning but remain underexplored in Abstract Visual Reasoning (AVR), which demands higher-order reasoning to identify abstract rules beyond simple perception. Existing AVR benchmarks focus on single-step reasoning, emphasizing the end result but neglecting the multi-stage nature of reasoning process. Past studies found MLLMs struggle with these benchmarks, but it doesn’t explain how they fail. To address this gap, we introduce MultiStAR, a Multi-Stage AVR benchmark based on RAVEN, to assess reasoning across varying levels of complexity. Additionally, existing metrics like accuracy only focus on the final outcomes while do not account for the correctness of intermediate steps. Therefore, we propose a novel metric, MSEval, which considers the correctness of intermediate steps in addition to the final outcomes. We conduct comprehensive experiments on MultiStAR using 17 representative close-source and open-source MLLMs. The results reveal that while existing MLLMs perform adequately on basic perception tasks, they continue to face challenges in more complex rule detection stages. The dataset and code are available at https://github.com/YanbeiJiang/MultiStAR
@@
Figure 1: Left: RAVEN puzzle. The correct answer is 1. Right: Direct Answer subtask, where questions are independent for each configuration; Logical Chain subtask, where information from previous stages is used to assist in answering the current stage. All questions here focus on the concept of Number.
@@ 1 Introduction
Multimodal Large Language Models (MLLMs) demonstrate proficiency in addressing a wide array of visual-text inquiries and show strong multimodal understanding ability in tasks such as visual question answering Goyal et al. (2017); Marino et al. (2019); Ding et al. (2023), image captioning Saito et al. (2023); Vinyals et al. (2015); Jiang et al. (2024a), and visual grounding He et al. (2024); Deng et al. (2021). These tasks focus on evaluating the models’ capability to understand real-world or domain-specific knowledge. However, Abstract Visual Reasoning (AVR) presents a different challenge, focusing on a model’s ability to identify and reason through abstract patterns, relationships, and rules. A well-known example of AVR tasks is RAVEN Raven (2003); Zhang et al. (2019), as shown in the left half of Figure 1. The solver needs to select the correct panel from a answer set to complete a 3x3 problem matrix by deducing the visual rules governing the grid’s arrangement. For instance, by analyzing the colors of each panel, one might observe the color remains consistent across each row. Unlike other multimodal tasks in real-world scenarios, AVR focuses on reasoning about arbitrary visual elements, serving as a robust benchmark for evaluating the zero-shot reasoning capabilities of MLLMs in visual contexts Mańdziuk and Żychowski (2019); Santoro et al. (2018).
@@
Previous works have consistently shown that AVR tasks pose challenges for MLLMs in zero-shot inference settings. Despite recent advancements like Chain-of-Thought prompting Ahrabian et al. (2024); Gendron et al. (2024) and the inclusion of oracle captions Zhang et al. (2024), models continue to perform at near-random levels on these tasks. The AVR datasets used commonly in these evaluations like RAVEN primarily focus on single-step end-to-end reasoning (i.e., giving the models the questions and asking them to derive the final answer; left half of Figure 1) Santoro et al. (2018); Nie et al. (2020); Cao et al. (2024). However, this design deviates from the human reasoning process, which often involves sequential steps: starting with single-panel perception, progressing to panel comparisons, and finally deducing the underlying rules before solving the puzzle. Previous datasets often omit these intermediate stages, and as such cannot evaluate MLLM’s step-by-step reasoning capabilities and identify where they struggle within the reasoning process. This highlights the need for benchmarks that assess intermediate perception and reasoning processes. Additionally, existing evaluation metrics like accuracy measure only the performance of the final answer without considering the correctness of intermediate steps., and so they are unable to assess partial success.
@@
To address these limitations, we introduce MultiStAR, a Multi-Stage Abstract Visual Reasoning dataset based on RAVEN, to evaluate MLLMs on the intermediate steps in the reasoning process. As shown in the right half of Figure 1, the dataset is divided into two sub-tasks, each focusing on different aspects of reasoning. The first sub-task, referred to as Direct Answer, evaluates model performance at varying levels of complexity to assess perception and reasoning abilities at each individual step. Using template-based methods, we generate questions based on RAVEN, ranging from basic object recognition to advanced comparison, pattern recognition, and rule inference. This approach ensures comprehensive coverage of reasoning patterns. The second sub-task called Logical Chain, emphasizes how models measure and maintain logical correlations across reasoning steps. Using puzzles from the RAVEN as the final question, we decompose the reasoning process into a sequence of subproblems in a bottom-up manner. Each stage in this chain links the current reasoning task to its dependent subproblems, requiring the model to combine current information with outputs from previous stages. To assess the correctness of intermediate steps, we introduce a novel metric, MSEval, which provides a more fine-grained assessment of the model’s reasoning process for the logical chain task. MSEval uses the correct answer probabilities at each stage to compute the joint probability across the reasoning process. This approach considers both the correctness of the current stage and all dependent intermediate steps.
@@
In summary, our contributions are: 1) we introduce the MultiStAR benchmark, designed to evaluate models across different stages of reasoning through two subtasks, allowing for a more granular analysis of their performance throughout the reasoning process; 2) we present a novel metric that incorporates the correctness of the current stage as well as the accuracy of its dependent intermediate steps; and 3) we perform extensive experiments on a wide range of state-of-the-art MLLMs, our results indicate that current VLLMs still lack the ability to navigate multi-stage logical dependencies effectively.
Figure 2: Left: Direct Answer subtask, showcasing six configurations along with their corresponding examples. Right: Logical Chain task, presenting a partial view of the logical chain (see full chain and the chain designing rationale in Appendix A.3). Examples are provided for one specific path in the chain.
@@ 2 Related Work
Visual reasoning benchmarks have evolved to assess the capacity of AI models on various tasks including compositional Johnson et al. (2017), commonsense Gao et al. (2022); Li et al. (2024), scientific Hiippala et al. (2021); Saikh et al. (2022); Yue et al. (2024), and abstract visual reasoning. Both commonsense and scientific reasoning tasks require real-world knowledge and a prior understanding of specific domains. Abstract Visual Reasoning (AVR) is the main focus of this work, and it primarily involves a classification task where models select an answer from a fixed set of choices based on abstract patterns and rules Mańdziuk and Żychowski (2019); Zhang et al. (2019); Santoro et al. (2018); Nie et al. (2020). A few other AVR benchmarks address generative tasks, where models are tasked with recreating elements that fit within a given visual sequence, introducing additional complexity by evaluating a model’s creative reasoning capabilities Chollet (2019); Moskvichev et al. (2023). The most similar benchmark to ours is MARVEL Jiang et al. (2024b), which targets AVR tasks and extends reasoning diversity with six core patterns across geometric and abstract shapes. It also includes basic perception questions to assess visual comprehension. However, MARVEL is still limited in its capacity to analyze intermediate reasoning steps. Table 1 shows the key statistics and features comparison of major multimodal reasoning datasets alongside our proposed MultiStAR benchmark.
@@ 4 Results
In the Direct Answer subtask, we evaluate 17 representative MLLMs in zero-shot setting, including both open-source and close-source models. For open-source models, we have vanilla pre-trained (i.e., without instruction-tuning) and instruction-tuned models. Additionally, we consider a range of model sizes to ensure a comprehensive evaluation. The model settings and the input prompt details are shown in Appendix A.4.1.
@@ 4.1 Result Analysis
Table 2 compares the performance of various MLLMs on the Direct Answer task. The two close-source models, GPT-4o and Gemini-1.5-pro, outperform others in 1P-B, 1R and 2R stages. GPT-4o achieves an impressive 88.07% accuracy for basic object-oriented questions within a single panel, highlighting its strong capability to recognize simple visual patterns. Gemini achieves the best results in rule deduction tasks for both one-row and two-row configurations, indicating its superior ability to process more complex visual inputs and perform logical reasoning effectively. Among open-source models, pre-trained models generally perform worse than instruction-tuned ones, which is no surprise as the pre-trained models are unlikely to be able follow the prompt/instructions. Qwen2-VL-Instruct-72B achieves the best performance on cross-panel comparison tasks. Interestingly, we also see that it has a very strong performance in the original RAVEN task (“Final”), and we suspect there is probably contamination since it struggles with the more complex intermediate reasoning tasks (“1R” and “2R”). We also think Idefics2-8B maybe be contaminated, again for the same reason.
We performed an error analysis (see Appendix 11) and identified that the models mainly face challenges with deep reasoning errors. Additionally, the decoder-only models (Qwen2-VL, InternVL2 and NVLM-D) tend to perform much better than encoder-decoder models (InstructBLIP and LLaVA-v1.5). This suggests that decoder-only architectures may be more effective for step-by-step reasoning to some extent. However, it’s important to note that the decoder-only models are newer and the performance gap may be attributed to other factors such as larger or more diverse training data or better training objectives.
Interestingly, as the questions become increasingly complex and require deeper reasoning, a noticeable decline in performance is observed across all models, gradually approaching the random baseline. While models demonstrate strong performance on basic perception tasks, they struggle significantly with deeper reasoning challenges. In contrast, human performance remains stable at above 60% with increasing complexity. This highlights the substantial gap between model and human performance, emphasizing the limitations of current MLLMs in understanding and reasoning at a level comparable to humans. Another finding is that performance generally improves with larger model sizes — the size of the base language model (see Appendix A.9.1 for detailed visual analysis). This suggests that these VLLMs might be primarily relying on its language model for tackling these AVR tasks.
@@
Table 2: The answer accuracy of MLLMs for the Direct Answer subtask. The best results are highlighted in bold. denotes contaminated models that may have been trained with RAVEN. means the number is taken from from Zhang et al. (2019).
@@
Table 3: The Accuracy (Acc) and MSEval scores for the Logical Chain task. denotes contaminated models that may have been trained with RAVEN. The highest accuracy are highlighted in bold. The highest MSEval are highlighted in underline. w/o: without prior, w: with prior. See Appendix A.4.2 for results of more models.
@@
Table 3 shows the performance of MLLMs on each stage of the Logical Chain task. Results are reported as accuracy and MSEval scores. Among all models, Gemini-1.5-pro achieves the best performance on the first four stages. Among the open-source models, Qwen2-VL-72B outperforms others in both accuracy and MSEval, although as mentioned earlier we think the model has been trained with RAVEN and we should probably interpret these results with caution (same for Idefics2-8B). When comparing results with and without prior information, all models show better performance when prior information is available in both metrics, highlighting their capacity to benefit from step-by-step reasoning, even when some generated previous answers might be incorrect. For visual representation of the percentage increase, see to Appendix A.7.
Interestingly, for the final stage involving the RAVEN puzzle, prior information appears to provide limited utility for most models (exception: Gemini). This aligns with previous findings that chain-of-thought reasoning models struggle to solve RAVEN puzzles Ahrabian et al. (2024); Gendron et al. (2024). However, MSEval scores tell a different story. Almost all models (exception: Qwen2-VL-72B) see an improvement in MSEval even though there is no change (or even a drop) in accuracy when prior information is incorporated. Since MSEval evaluates correctness across intermediate and current stages, it reveals that even though prior information might not help MLLMs solve the final stage question, it helps them with the intermediate tasks.
Another finding is the MSEval score for Qwen2-VL-72B declines when provided with prior information, which is not consistent with accuracy. This indicates its weakness in addressing intermediate stages that were likely not part of its pre-training. Despite errors in earlier stages, the model still performs well on the final question, suggesting it relies on memorizing patterns from the final stage rather than demonstrating a strong understanding of the logical reasoning behind the task. This highlights a critical limitation in current MLLMs, while they may achieve impressive results in isolated cases, their ability to generalize and reason through multi-stage logical dependencies remains inadequate. To further verify the MSEval’s effectiveness, we also conduct qualitative analysis in section 6.
@@ 5 Discussion
From Figure 5, the “number” attribute is the easiest to recognize in higher level configurations (2P, 1R, 2R), while “position” is the most easily identified in low-level, single-panel settings. Some attributes achieve accuracy above 90%, indicating that the models exhibit strong counting and spatial reasoning capabilities. However, they struggle with attributes like “color” and “size”, particularly in high-level configurations.
Figure 5: Breakdown analysis of five attributes for Gemeni-1.5-pro and Qwen2-VL-72B on the Direct Answer task. See Appendix A.9.2 for more models.
Given the ground truth (i.e., correct) answer for intermediate steps, how does it influence the final results?
Table 4 highlights that the correctness of the prior information is important. For example, the 1R stage benefits significantly from the insights about each panel and intra-panel comparisons. The 2R stage also sees substantial gains, as it mainly relies on double-checking information from the 1R stage without requiring additional changes in most cases. However, the final stage experiences a negative impact despite the inclusion of correct rules. This may be attributed to the complexity of the visual input, which contains numerous objects, making it challenging for the model to effectively apply the given rules. And for the Qwen2-VL-72B model, its tendency to memorize patterns might turn these ground truths into noise.
[…]
@@
[…]
Table 4: The accuracy with incorporating ground truth information at each stage for Logical Chain task.
@@
Injecting prior information into prompts significantly increases their length (see Appendix A.1 for details), making it more challenging for models to focus on critical details. To address this issue, we proposed two methods: (1) adding HTML tags to structure the prompt by separating prior information, background, and questions, enabling the model to clearly distinguish each part, and (2) formatting the prompt as a PDF document with distinct sections and titles. Table 6 demonstrates that HTML tagging provides some benefits, but not so much for the PDF approach (see Appendix A.9.3 for further details and examples of the conversion methods).
[…]
@@
[…]
Table 6: The accuracy of three prompting techniques for prior information for Logical Chain task. Vanilla: Pure Text, Struct.: Structure (HTML), Doc.: Document.